OpenStudy (anonymous):

what is the length of a rectangle with width 10in and area 45in squared

OpenStudy (anonymous):

width*length=area. 10*x=45 x=45/10=4.5

OpenStudy (anonymous):

and that with it squared ?

OpenStudy (anonymous):

area is always in the units of in squared. dont pay any attention to the units. you're doing the math, so deal with the numbers. if the area was 45 squared, which is 2025, what would be a different story.

OpenStudy (anonymous):

im confused ?

OpenStudy (anonymous):

just know that you don't need to know about the inches squared part.

OpenStudy (anonymous):

ok thankyou!!!

OpenStudy (anonymous):

no problem